Operations Report – May 16th

Crew 212 Operations Report 16-MAY-2019

SOL: 11

Name of person filing report: Paolo Guardabasso

Non-nominal systems: None.

Notes on non-nominal systems: Nothing to report.

Generator (hours run last night): [From Director]
On (last night): 6:15 PM
Off (this morning): [From Director]

Solar— SOC 84% at 7:40 PM

Diesel Reading – 82%

Propane Reading – 45%

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 200 gallons

Axillary to Static tank transfer – no

Gallons transferred: No

Water in GreenHab – 113 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 143754.6

Toilet tank emptied: No

Deimos rover used: no (out of order)
Hours:
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes [USED BY DIRECTOR]
Hours: [From Director]
Beginning charge: 100%
Ending charge: 71%
Currently charging: Yes

Opportunity rover used: Yes [USED BY DIRECTOR]
Hours: [From Director]
Beginning charge: 76%
Ending charge: 60%
Currently charging: Yes

Curiosity rover used: No
Hours: 96.6h
Beginning charge:
Ending charge:
Currently charging: Yes

Notes on rovers: [From Director] Sojo’s battery bank was cleaned and batteries refilled with distilled water. Only one cell was critically low (below the metal) but it was still wet. 1.5 gallons of distilled water used. No connections were loose and no battery caps were trapped under the security bar. The thermistor on the Rangers need to go to get serviced. All the batteries on Oppi were cleaned and the batteries will be filled tomorrow. Oppi needs the brakes serviced, as they have failed and run out of brake fluid twice this season.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? Nothing to report.

CrewCar used and why, where? Off Campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: after EVA#11, we had a debriefing and we worked on our experiments. The high winds have slightly damaged the tunnel covers in a few points.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: the new pump is working fine, but I just wanted a clarification: is it normal that the pump works in a pulsed mode (oscillating ON/OFF) when a low quantity of water is required?